Shift report 20011109 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Nov 09 2001 - 16:02:58 EST

  • Next message: Stephen J. Sanders: "tenure track faculty position"

    ------------------------------------------------------------
    Date: Fri Nov  9 16:02:57 2001, Shift: 08:00-16:00
    Supervisor: TST, with: 
    Log book#10 pages 118 - 119
    
    Shift Summary:
    ------------------------------------------------------------
    1:  Store from previous run continued to about 11.15,
    with falling intensity.  Last part of store used for
    beam studies - stopped data taking.  Injection attempts
    around end of shift.
    
    2:  Run   Trig1  Trig6    
        5737   380k   120k   MRS@90,4B  FFS@20,1/15B
        5738   148k    46k   MRS@90,0   FFS@20,1/10B
        5739                 DVM pedestal run
        5740                 Pedestal run
        5731                 TPC pedestal run
    
    3:  No new detector problems.  All DC anodes still off.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Nov 09 2001 - 16:03:53 EST